A disk rolls smoothly down an incline. (a) If α = 250, calculate the translational acceleration of the center of mass of the disk. (I = 1/2 MR2); (b) Analyze forces acting on the disk. What forces cause the translational motion and the rotational motion of the disk?


1
Expert's answer
2021-07-01T04:56:38-0400

a)


"a=g\sin{25}\\a=9.8\sin{25}=4.1\frac{m}{s^2}"

b) Weight and frictional force.